A painter can paint a racquetball court in 8 hours. His assistant needs an additional 2 hours to paint the same court working by himself. How long will it take both of them working together to paint the court. Thank You
Answer by scott8148(6628)

each one paints a FRACTION of the court , based on their individual rates
___ the fractions sum to 1 (the whole court)

let x="combined time"

(x/8) + [x/(8+2)] = 1 ___ (x/8) + (x/10) = 1

multiplying by 40 (LCD) ___ 5x + 4x = 40 ___ 9x = 40

dividing by 9 ___ x = 40/9

4 hr 26 min 40 sec